# Breaking Barriers or Building Dependency? Exploring Team-LLM Collaboration in AI-infused Classroom Debate

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in collaborative and social computing
@@ 1. Introduction
Researchers are increasingly interested in facilitating human-AI collaboration in complex learning environments (Amershi et al., 2019; Cai et al., 2019), recognizing AI’s potential to transform traditional learning approaches (Pataranutaporn et al., 2021). Therefore, employing AI in collaborative learning is becoming a significant area of focus. In this domain, a holistic perspective on human-AI partnering is essential, rather than focusing solely on either human or AI capabilities (Huang et al., 2019). Human-AI collaborative learning requires three key characteristics: ”mutual understanding,” ”mutual benefits,” and ”mutual growth” (Huang et al., 2019). While the optimal combination of human expertise and AI capability promises great potential, further exploration is urgently needed to fully comprehend and leverage these interactions (Islam et al., 2023).
@@
Our study aims to gain contextualized insights into human-AI hybrid collaborative learning by focusing specifically on AI-assisted classroom debates. Unlike informal debates or general discussions, formal debates are characterized by clear rules and structures, emphasizing the construction of logically rigorous arguments and strategic rebuttals (Massey and Meegan, 2015). We chose the classroom debate setting since it represents a high-paced, highly interactive learning activity involving intensive collaboration, communication, and competition within a limited time. This contrasts with the more commonly studied scenarios where human-AI interactions are analytical and relatively less intensive (Zheng et al., 2024). We believe that examining the dynamics of AI-infused classroom debates can result in unique and meaningful insights that will inform future research on human-AI hybrid collaborative learning across similar contexts such as classroom peer reviews (Mangelsdorf, 1992), Socratic seminars (Tredway, 1995), collaborative augmentation (Chinn and Clark, 2013) and problem-based learning (Wood, 2003).
@@
To do so, we organized 3 rounds of five-to-five classroom debates involving 22 students, aiming to understand how learners interact with AI and how these interactions impact their learning process and team collaboration. Each debate centered on a different theme regarding the course Design History, allowing us to observe how participants use AI support in their arguments. We examined learners’ emerging ways of interacting with an LLMs-based interface and how these interactions influenced their performance and team dynamics. During the debates, students were required to present their viewpoints and respond to opponents’ queries with support from AI-provided examples and perspectives. The entire activity was recorded for analysis. In addition, we conducted in-depth interviews with the 22 participants to explore their views of AI usage during the debates and how this usage influenced their learning behaviors and learning experiences. Analyzing these data provided us with a deeper understanding of potential benefits, as well as the potential challenges of human-AI collaborative learning in this context. Namely, our investigation encompassed two key research questions:
@@
By analyzing these facets, we aim to discern the impacts of integrating AI into educational debate scenarios. Additionally, the study probed how teams acquired information, formed consensus, and achieved common goals within collaborations and how AI-infused technology served as a mediator to facilitate knowledge sharing. Our findings indicate that AI tools can facilitate the debate and collaborative learning experience by enabling adaptive, tailored responses, offering diverse perspectives, reducing collaborative pressure, and accelerating information retrieval. However, we also observed potential risks, such as cognitive dependency, the undermining of individual creativity, and the challenges of shifting communication expectations. By addressing the authentic classroom debate setting, our research has uncovered contextual insights that may further future research and contribute to designing more effective and considerate human-AI collaborative learning systems that aim to meet broader educational contexts.
@@ 2.1.1. Traditional AI and Technology in Education
In the field of education, traditional AI is often applied in areas such as learning analytics (Zambrano and Baker, 2024; Dang et al., 2024; Nazaretsky et al., 2022), educational data mining (Borchers et al., 2024a; Castro et al., 2023), intelligent tutoring systems (ITS) (Weber et al., 2024; Zhou et al., 2020; Tang et al., 2023; Zhang et al., 2024; Borchers et al., 2024b; Booth et al., 2024), and learner profiling (Mejia-Domenzain et al., 2023). In these studies, AI is primarily used to analyze vast educational data, generating patterns and insights that help educators and learners reflect on the teaching and learning processes. These systems often provide post-hoc analysis to guide future educational activities or offer evidence-based insights for educational administrators. Some other works are focusing on interactive technologies, such as robots (Elmimouni et al., 2024), AR (Eisenberg et al., 2022; Camara Olim et al., 2023; Zhou et al., 2024; Jin et al., 2023; Caetano et al., 2023) and VR (Son et al., 2024; Jin et al., 2022), cloud-based and client-side applications (Underwood et al., 2023; Chatzidaki et al., 2024; Zhang et al., 2023a; Núñez et al., 2023; Ngoon et al., 2024), dashboards (Lee-Cultura et al., 2023; Fernandez-Nieto et al., 2024; Echeverria et al., 2024; Buvari et al., 2023; Karademir et al., 2022; Gauthier et al., 2022), spatial applications (Martinez-Maldonado et al., 2022; Fernandez-Nieto et al., 2022), and documentation tools (Tan and Subramonyam, 2024; Sterman et al., 2023), exploring how these technologies enhance teaching and learning activities. For example, the study by Elmimouni et al. investigated how telepresence robots both enable and limit classroom accessibility (Elmimouni et al., 2024). These studies encompass both formal learning and informal learning, providing rich design cases that show how various media can enhance subject-specific learning across different contexts.
@@ 2.1.3. LLMs-infused Multi-learner Collaborative Learning
However, much of the aforementioned research (Ma et al., 2024a; Zhang et al., 2024; Göldi et al., 2024; Hänel et al., 2024) focuses on the interaction between LLM systems and individual users, with relatively little attention given to multi-learner collaborative learning. A few examples explore LLMs in group conversation (Liu et al., 2024a), group brainstorming (Yu-Han and Chun-Ching, 2023), teamwork (Zhang et al., 2023b), group decision making (Chiang et al., 2024), and classroom simulation (Liu et al., 2024b). For example, Liu et al., explored the role of LLM agents in children’s collaborative learning, showing that these agents can effectively moderate discussions and foster creative thinking (Liu et al., 2024a). Additionally, Chiang et al. introduced an LLM-powered devil’s advocate, which enables teams to receive opposing viewpoints during the decision-making process, which fosters critical thinking and enhances the overall quality of decisions (Chiang et al., 2024).
While these collaborative activities leverage LLMs to facilitate multi-learner interaction, the scenarios explored in most studies generally do not emphasize strict time constraints and knowledge grasp and application. In contrast, classroom formal debate, as a high-paced and time-sensitive form of collaborative learning, focuses more on efficient teamwork and fast group learning than the aforementioned studies. In high-intensity scenarios, the emerging Team-LLM interaction modes exhibit notable nuances compared to low-intensity, unrestricted environments. These unique nuances of the debate offer valuable insights into the potential to use AI to support such dynamic learning settings.
@@ 2.2. Classroom Debate as Collaborative Learning
Debate is a structured dialogue where participants present, support, and refute viewpoints on a specific topic. It fosters a deeper understanding of complex issues and develops critical thinking, communication, and teamwork skills through logical arguments and evidence. Debate, as an educational strategy, promotes active learning by guiding learners from lower-level thinking to higher-order thinking, such as analysis and evaluation (Vo and Morris, 2006). It also fosters critical thinking by encouraging the exploration of diverse perspectives, improving understanding, and enhancing skills in communication and teamwork. Roy’s study found that classroom debates are more effective than traditional lectures in developing higher-order thinking skills (Roy and Macchiette, 2005). Classroom debates can be regarded as a form of collaborative learning, with related research showing that debates have a positive impact on learners’ communication and thinking skills. Brown’s study indicates that in real classroom environments, debates, as a form of collaborative learning, allow students to actively participate in discussions. This participation fosters interaction among students and strengthens the collaborative relationship between teachers and students as they work together toward a common goal. These are all essential skills in collaborative learning (Brown, 2015). In the communication process, Berdine points out that learners must temporarily set aside their biases. This practice of perspective-taking helps them become more open to different viewpoints, enhancing their ability to collaborate and communicate effectively within a team (Berdine, 1987). Additionally, Gregory et al. highlight that debates require both the audience and learners to evaluate different choices (Gregory and Holloway, 2005). Therefore, it aligns with the requirements for developing thinking skills by advancing through the levels of Bloom’s Taxonomy (Kennedy, 2009). Learners can actively analyze, discuss, and apply content meaningfully rather than passively absorbing information. This active logical thinking process allows for more effective learning (Bonwell and Eison, 1991), thereby improving their problem-solving skills. Therefore, Brown points out that learners need to evaluate and construct arguments during the debate process. This activity not only improves learners’ ability to solve complex problems but also enhances the development of their logical thinking skills (Brown, 2015). Furthermore, formal debates, as structured and regulated activities, are beneficial to enhance participants’ rhetorical skills and their adherence to the rules of decorum (Vo and Morris, 2006). In summary, the learning objectives of formal debate include gaining a deeper understanding of complex issues, enhancing critical thinking, improving communication and teamwork skills, fostering higher-order thinking such as analysis and evaluation, mastering rhetoric, and adhering to the rules of decorum.
@@
The integration of group learning theories and AI-supported debates provides a robust framework for enhancing collaborative learning. Vygotsky’s Zone of Proximal Development (ZPD) (Vygotsky, 1978) underscores the importance of scaffolding, such as adult guidance, peer collaboration, or technological support, in enabling learners to surpass their independent capabilities. Sociocultural theory, as articulated by Mercer (Mercer, 2002, 2007), highlights the pivotal role of language in linking social interaction with cognitive development, where exploratory talk serves as a critical medium for knowledge construction. Furthermore, Intersubjectivity, as described by Rogoff and Wertsch (Rogoff, 1990; Wertsch, 1991b, a), emphasizes the importance of shared understanding in collaborative tasks. The concept of Sociol-cognitive Conflict, proposed by Doise and Mugny (Doise, 1984), identifies cognitive adjustment through conflict and negotiation as a mechanism for deep learning. Finally, Exploratory Talk, as defined by Mercer (Mercer, 2002), is a high-quality dialogic form that enhances reasoning and collaborative skills through critical negotiation and reasoned analysis.
@@
Debate is a language-mediated thought process that can seamlessly integrate with LLMs, which also function through language. As LLM technology advances, it shows great potential in enhancing language-based learning activities across both formal and informal educational settings, such as creative writing (Göldi et al., 2024; Mejia-Domenzain et al., 2024; Tu et al., 2024), public speaking (Park and Choi, 2023; Follmer et al., 2023), second language acquisition (Zhang et al., 2024; Kim et al., 2024; Higasa et al., 2024; Ma et al., 2024a), and reading (Hänel et al., 2024; Xu et al., 2023). However, LLM-based systems have not yet been widely applied for debate assistance, with only a few studies exploring how LLMs can support learners during the debate preparation phase (Xu et al., 2023; Guo et al., 2023; Li et al., 2024). Nevertheless, considering the advantages of LLMs in processing and generating language, we believe that applying LLM technology in classroom debates can bring significant benefits to a wide range of educational and learning contexts and collaborative learning. For example, in-classroom peer reviews are time-sensitive (during class sessions) and require participants to engage in comprehensive and critical evaluations of others’ writing. This process demands thorough consideration and necessitates providing academically rigorous feedback (Mangelsdorf, 1992). Similarly, Socratic seminars emphasize the cultivation of common values and free inquiry. Participants in a debate are expected to establish a foundation of shared principles (common values), which fosters a unified stance. At the same time, the spirit of free inquiry encourages participants to engage in critical thinking and dialectical reasoning. This involves collaboratively brainstorming arguments, exploring diverse cases, and anticipating counterarguments that might arise during the debate. Likewise, other learning contexts such as collaborative augmentation (Chinn and Clark, 2013) and problem-based learning (Wood, 2003) also align seamlessly with the objectives of this research on AI-infused classroom debates, which aim to promote critical analysis, cooperative argumentation, and the development of shared understanding in time-sensitive educational settings, which are areas that traditional research on collaborative learning has paid relatively little attention to. In human-AI interaction research, classroom debates, as a structured dialogue and educational strategy, offer a unique context for human-AI collaborative learning. The fast pace, high intensity, and need for interpersonal collaboration in debates highlight their educational value. This study aims to address a gap by exploring AI’s role in enhancing collaborative learning.
@@ 3.1. Study Setup
The study was conducted in the Design History course at Anonymous University, where all the participants were enrolled. The course adopted an active learning approach, and classroom debates had been originally planned to facilitate a cooperative and learner-driven environment (Slavin, 1996). Initially, potential participants who were students enrolled in the Design History course were contacted to verify their willingness to engage in the study. As a result, 22 students participated, composed of 11 females and 11 males. As shown in Fig. 2, the distribution of their academic years was 1 freshman, 20 sophomores, and 1 senior. Participants rated their level of previous experience with using Large Language Models (LLMs) at 3.50 (SD = 0.72) on a 1 to 5 scale (no experience to very experienced). Detailed questionnaire and survey results are shown in Appendix A.
@@
Three rounds of five-to-five debates were conducted over four weeks, with all participants present in each round. Each round of debates took approximately 1.5 hours. The participants were randomly assigned to either the debate’s affirmative or negative side and each side consisted of five team members. As the audience, other participants were tasked with observing the debates and voting at the beginning and end, and audience members also helped with timekeeping. The teams were able to use their digital devices during the debate. The rules stipulate that only one laptop can be accessed and use ChatGPT 3.5 in a group with no additional restrictions. By restricting AI usage to a single device, we encourage participants to develop their collaborative strategies autonomously rather than solely using AI. Rich data were gathered through classroom observations and interviews, aimed at deepening the understanding of participants’ metacognitive processes (i.e., assessment and monitoring of their cognitive operations in the AI-assisted debates (Mahdavi, 2014)) and their experience of teamwork dynamics. The debate process was designed to fit the classroom environment, structured into the following stages (see Fig. 3):
@@
    Team Summary: Each team has 5 minutes for their concluding remarks.
@@
This study has been reviewed and approved by the authors’ institution (IRB No.: 20240274). All the participants gave informed consent, and the research data was anonymized and handled carefully.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06journalyear: 2025copyright: acmlicensedconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713853isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Empirical studies in collaborative and social computing
@@ 1. INTRODUCTION
Researchers are increasingly interested in facilitating human-AI collaboration in complex learning scenarios (Amershi et al., 2019; Cai et al., 2019), recognizing AI’s potential to transform traditional learning approaches (Pataranutaporn et al., 2021). Therefore, employing AI in collaborative learning is becoming a significant area of focus. In this domain, a holistic perspective on human-AI partnering is essential, rather than focusing solely on either human or AI capabilities (Huang et al., 2019). Human-AI collaborative learning requires three key characteristics: ”mutual understanding,” ”mutual benefits,” and ”mutual growth” (Huang et al., 2019). While the optimal combination of human expertise and AI capability promises great potential, further exploration is urgently needed to fully comprehend and leverage these interactions (Islam et al., 2023).
@@
Our study aims to gain contextualized insights into human-AI hybrid collaborative learning by focusing specifically on AI-infused classroom debates. Unlike informal debates or general discussions, formal debates are characterized by clear rules and structures, emphasizing the construction of logically rigorous arguments and strategic rebuttals (Massey and Meegan, 2015). We chose the classroom debate setting since it represents a high-paced, highly interactive learning activity involving intensive collaboration, communication, and competition within a limited time. This contrasts with the more commonly studied scenarios where human-AI interactions are analytical and relatively less intensive (Zheng et al., 2024). We believe that examining the dynamics of AI-infused classroom debates can result in unique and meaningful insights that will inform future research on human-AI hybrid collaborative learning across similar contexts, such as classroom peer reviews (Mangelsdorf, 1992), Socratic seminars (Tredway, 1995), collaborative augmentation (Chinn and Clark, 2013) and problem-based learning (Wood, 2003).
@@
To do so, we organized 3 rounds of five-to-five classroom debates involving 22 students to understand how learners interact with AI and how these interactions impact their learning process and team collaboration. Each debate centered on a different theme regarding the course Design History, allowing us to observe how participants use AI support in their arguments. We examined learners’ emerging ways of interacting with an LLMs-based interface and how these interactions influenced their performance and team dynamics. During the debates, students were required to present their viewpoints and respond to opponents’ queries with support from AI-provided examples and perspectives. The entire activity was recorded for analysis. In addition, we conducted in-depth interviews with the 22 participants to explore their views of AI usage during the debates and how this usage influenced their learning behaviors and learning experiences. Analyzing these data provided us with a deeper understanding of the potential benefits and challenges of human-AI collaborative learning in this context. Namely, our investigation encompassed two key research questions:
@@
By analyzing these facets, we aim to reveal the impacts of integrating AI into educational debate scenarios. Additionally, the study probed how teams acquired information, formed consensus, and achieved common goals within collaborations and how AI-infused applications served as a mediator to facilitate knowledge sharing. Our findings indicate that AI tools can facilitate the debate and collaborative learning experience by enabling adaptive, tailored responses, offering diverse perspectives, reducing collaborative pressure, and accelerating information retrieval. However, we also observed potential risks, such as cognitive dependency, the undermining of individual creativity, and the challenges of shifting communication expectations. Our research is grounded in an authentic classroom debate environment, offering contextual insights that may guide future research and contribute to designing more effective and considerate human-AI collaborative learning systems for broader educational contexts.
@@ 2.1.1. Traditional AI and Technology in Education
In the field of education, traditional AI is often applied in areas such as learning analytics (Zambrano and Baker, 2024; Dang et al., 2024; Nazaretsky et al., 2022), educational data mining (Borchers et al., 2024a; Castro et al., 2023), intelligent tutoring systems (ITS) (Weber et al., 2024; Zhou et al., 2020; Tang et al., 2023; Zhang et al., 2024; Borchers et al., 2024b; Booth et al., 2024), and learner profiling (Mejia-Domenzain et al., 2023). In these studies, AI is primarily used to analyze vast educational data, generating patterns and insights that help educators and learners reflect on the teaching and learning processes. These systems often provide post-hoc analysis to guide future educational activities or offer evidence-based insights for educational administrators. Some other works are focusing on interactive technologies, such as robots (Elmimouni et al., 2024), AR (Eisenberg et al., 2022; Camara Olim et al., 2023; Zhou et al., 2024; Jin et al., 2023; Caetano et al., 2023) and VR (Son et al., 2024; Jin et al., 2022), cloud-based and client-side applications (Underwood et al., 2023; Chatzidaki et al., 2024; Zhang et al., 2023a; Núñez et al., 2023; Ngoon et al., 2024), dashboards (Lee-Cultura et al., 2023; Fernandez-Nieto et al., 2024; Echeverria et al., 2024; Buvari et al., 2023; Karademir et al., 2022; Gauthier et al., 2022), spatial applications (Martinez-Maldonado et al., 2022; Fernandez-Nieto et al., 2022), and documentation tools (Tan and Subramonyam, 2024; Sterman et al., 2023), exploring how these technologies enhance teaching and learning activities. For example, the study by Elmimouni et al. investigated how telepresence robots both enable and limit classroom accessibility (Elmimouni et al., 2024). These studies encompass both formal and informal learning, providing rich design cases showing how various media can enhance subject-specific learning across different contexts.
@@ 2.1.3. LLMs-infused Multi-learner Collaborative Learning
However, much of the aforementioned research (Ma et al., 2024a; Zhang et al., 2024; Göldi et al., 2024; Hänel et al., 2024) focuses on the interaction between LLM systems and individual users, with relatively little attention given to multi-learner collaborative learning. A few examples explore LLMs in group conversation (Liu et al., 2024a), group brainstorming (Yu-Han and Chun-Ching, 2023), teamwork (Zhang et al., 2023b), group decision making (Chiang et al., 2024), and classroom simulation (Liu et al., 2024b). For example, Liu et al. explored the role of LLM agents in children’s collaborative learning, showing that these agents can effectively moderate discussions and foster creative thinking (Liu et al., 2024a). Additionally, Chiang et al. introduced an LLM-powered devil’s advocate, which enables teams to receive opposing viewpoints during the decision-making process, which fosters critical thinking and enhances the overall quality of decisions (Chiang et al., 2024).
While these collaborative activities leverage LLMs to facilitate multi-learner interaction, the scenarios explored in most studies generally do not emphasize strict time constraints and knowledge grasp and application. In contrast, as a high-paced and time-sensitive form of collaborative learning, classroom formal debate focuses more on efficient teamwork and fast group learning than the aforementioned studies. In high-intensity scenarios, the emerging Team-LLM interaction modes exhibit notable nuances compared to low-intensity, unrestricted environments. These unique nuances of the debate offer valuable insights into the potential to use AI to support such dynamic learning settings.
@@ 2.2. Classroom Debate as Collaborative Learning
Debate is a structured dialogue where participants present, support, and refute viewpoints on a specific topic. It fosters a deeper understanding of complex issues and develops critical thinking, communication, and teamwork skills through logical arguments and evidence. Debate, as an educational strategy, promotes active learning by guiding learners from lower-level thinking to higher-order thinking, such as analysis and evaluation (Vo and Morris, 2006). It also fosters critical thinking by encouraging the exploration of diverse perspectives, improving understanding, and enhancing skills in communication and teamwork. Roy’s study found that classroom debates are more effective than traditional lectures in developing higher-order thinking skills (Roy and Macchiette, 2005). Classroom debates can be regarded as a form of collaborative learning, with related research showing that debates positively impact learners’ communication and thinking skills. Brown’s study indicates that debates, as a form of collaborative learning, allow students to actively participate in discussions in real classroom environments. This participation fosters student interaction and strengthens the collaborative relationship between teachers and students as they work together toward a common goal. These are all essential skills in collaborative learning (Brown, 2015).
In the communication process, Berdine points out that learners must temporarily set aside their biases. This practice of perspective-taking helps them become more open to different viewpoints, enhancing their ability to collaborate and communicate effectively within a team (Berdine, 1987). Additionally, Gregory et al. highlight that debates require both the audience and learners to evaluate different choices (Gregory and Holloway, 2005). Therefore, it aligns with the requirements for developing thinking skills by advancing through the levels of Bloom’s Taxonomy (Kennedy, 2009). Learners can actively analyze, discuss, and apply content meaningfully rather than passively absorbing information. This active logical thinking process allows for more effective learning (Bonwell and Eison, 1991), thereby improving their problem-solving skills. Therefore, Brown points out that learners need to evaluate and construct arguments during the debate process. This activity not only improves learners’ ability to solve complex problems but also enhances the development of their logical thinking skills (Brown, 2015). Furthermore, formal debates, as structured and regulated activities, are beneficial to enhance participants’ rhetorical skills and their adherence to the rules of decorum (Vo and Morris, 2006). In summary, the learning objectives of formal debate include gaining a deeper understanding of complex issues, enhancing critical thinking, improving communication and teamwork skills, fostering higher-order thinking such as analysis and evaluation, mastering rhetoric, and adhering to the rules of decorum.
@@
Integrating group learning theories and AI-supported debates provides a robust framework for enhancing collaborative learning. Vygotsky’s Zone of Proximal Development (ZPD) (Barrs and Richmond, 2024) underscores the importance of scaffolding, such as adult guidance, peer collaboration, or technological support, in enabling learners to surpass their independent capabilities. Sociocultural theory, as articulated by Mercer (Mercer, 2002, 2007), highlights the pivotal role of language in linking social interaction with cognitive development, where exploratory talk serves as a critical medium for knowledge construction. Furthermore, Intersubjectivity, as described by Rogoff and Wertsch (Rogoff, 1990; Wertsch, 1991b, a), emphasizes the importance of shared understanding in collaborative tasks. The concept of Sociol-cognitive Conflict, proposed by Doise and Mugny (Doise, 1984), identifies cognitive adjustment through conflict and negotiation as a mechanism for deep learning. Finally, Exploratory Talk, as defined by Mercer (Mercer, 2002), is a high-quality dialogic form that enhances reasoning and collaborative skills through critical negotiation and reasoned analysis.
@@
Debate is a language-mediated thought process that can seamlessly integrate with LLMs, which also function through language. As LLM technology advances, it shows great potential in enhancing language-based learning activities across both formal and informal educational settings, such as creative writing (Göldi et al., 2024; Mejia-Domenzain et al., 2024; Tu et al., 2024), public speaking (Park and Choi, 2023; Follmer et al., 2023), second language acquisition (Zhang et al., 2024; Kim et al., 2024; Higasa et al., 2024; Ma et al., 2024a), and reading (Hänel et al., 2024; Xu et al., 2023). However, LLM-based systems have not yet been widely applied for debate assistance, with only a few studies exploring how LLMs can support learners during the debate preparation phase (Xu et al., 2023; Guo et al., 2023; Li et al., 2024). Nevertheless, considering the advantages of LLMs in processing and generating language, we believe that applying LLM technology in classroom debates can significantly benefit a wide range of educational and learning contexts and collaborative learning. For example, in-classroom peer reviews are time-sensitive (during class sessions) and require participants to engage in comprehensive and critical evaluations of others’ writing. This process demands thorough consideration and necessitates providing academically rigorous feedback (Mangelsdorf, 1992). Similarly, Socratic seminars emphasize the cultivation of common values and free inquiry. Participants in a debate are expected to establish a foundation of shared principles (common values), which fosters a unified stance. At the same time, the spirit of free inquiry encourages participants to engage in critical thinking and dialectical reasoning. This involves collaboratively brainstorming arguments, exploring diverse cases, and anticipating counterarguments that might arise during the debate. Likewise, other learning contexts such as collaborative augmentation (Chinn and Clark, 2013) and problem-based learning (Wood, 2003) also align seamlessly with the objectives of this research on AI-infused classroom debates, which aim to promote critical analysis, cooperative argumentation, and the development of shared understanding in time-sensitive educational settings, which are areas that traditional research on collaborative learning has paid relatively little attention to. In human-AI interaction research, classroom debates, as a structured dialogue and educational strategy, offer a unique context for human-AI collaborative learning. The fast pace, high intensity, and need for interpersonal collaboration in debates highlight their educational value. Our study aims to address a gap by exploring AI’s role in enhancing collaborative learning.
@@ 3.1. Study Setup
The study was conducted in the Design History course at the School of Design, Southern University of Science and Technology (SUSTech), where all the participants were enrolled. The course adopted an active learning approach, and classroom debates had been originally planned to facilitate a cooperative and learner-driven environment (Slavin, 1996). Initially, potential participants who were students enrolled in the Design History course were contacted to verify their willingness to engage in the study. As a result, 22 students participated, composed of 11 females and 11 males. As shown in Fig. 2, the distribution of their academic years was 1 freshman, 20 sophomores, and 1 senior. Participants rated their level of previous experience with using Large Language Models (LLMs) at 3.50 (SD = 0.72) on a 1 to 5 scale (no experience to very experienced). Detailed questionnaire and survey results are shown in Appendix A.
@@
Three rounds of five-to-five debates were conducted over four weeks, with all participants present in each round. Each round of debates took approximately 1.5 hours. The participants were randomly assigned to either the debate’s affirmative or negative side, and each side consisted of five team members. As the audience, other participants were tasked with observing the debates and voting at the beginning and end, and audience members also helped with timekeeping. The teams were able to use their digital devices during the debate. The rules stipulate that only one laptop can be accessed and use ChatGPT 3.5 in a group with no additional restrictions. By restricting AI usage to a single device, we encourage participants to develop their collaborative strategies autonomously rather than solely using AI. Rich data were gathered through classroom observations and interviews, aimed at deepening the understanding of participants’ metacognitive processes (i.e., assessment and monitoring of their cognitive operations in the AI-assisted debates (Mahdavi, 2014)) and their experience of teamwork dynamics. The debate process was designed to fit the classroom environment, structured into the following stages (see Fig. 3):
@@
    Team Summary: Each team has 5 minutes for concluding remarks.
@@
The authors’ institution reviewed and approved this study (IRB No.: 20240274). All the participants gave informed consent, and the research data was anonymized and used only for academic purposes.